OpenStudy (anonymous):

what is the answer of : nthroot 145

OpenStudy (anonymous):

? use the equation editor to write please...

OpenStudy (anonymous):

\[\sqrt[n]{145}=?\]

OpenStudy (anonymous):

here : \[\sqrt{145}\]

OpenStudy (anonymous):

im not that smart just so u know and its due tmrw and i have 3 exams :(

OpenStudy (anonymous):

can you use a calculator to answer?

OpenStudy (anonymous):

uumm ... it doesnt give me anything :(

OpenStudy (jdoe0001):

hmmmm what's the context for \(\huge \sqrt[n]{145}\quad ?\)

OpenStudy (anonymous):

it should be close to 12 because \[12^{2}=144 \Rightarrow \sqrt{144}=12\]

OpenStudy (anonymous):

yeah i know that but i need

OpenStudy (anonymous):

the other one

OpenStudy (anonymous):

http://web2.0calc.com/ try using this. just enter 145 and then click the square root button and then the equals button.